(x + 21)^2 = x^2 + 1869 X^2 + 441 = X^2 + 1869 x^2 + 441 - 441 = X^2 + 1869 - 441 x^2 = x^2 + 1428 Stuck here.

Algebra ->  Surface-area -> SOLUTION: Ref. Question 1048321: (x + 21)^2 = x^2 + 1869 X^2 + 441 = X^2 + 1869 x^2 + 441 - 441 = X^2 + 1869 - 441 x^2 = x^2 + 1428 Stuck here.      Log On


   



Question 1048398: Ref. Question 1048321:

(x + 21)^2 = x^2 + 1869
X^2 + 441 = X^2 + 1869
x^2 + 441 - 441 = X^2 + 1869 - 441
x^2 = x^2 + 1428
Stuck here.

Answer by Alan3354(69443)

(x + 21)^2 = x^2 + 1869
X^2 + 441 = X^2 + 1869 ******* (x+21)^2 = x^2 + 42x + 441
x^2 + 441 - 441 = X^2 + 1869 - 441
x^2 = x^2 + 1428